# 3DRot: 3D Rotation Augmentation for RGB-Based 3D Tasks

### v2

@@ Introduction
Experiments
Datasets and Metrics
We build on OmniNOCS (Krishnan et al. 2024) and focus on its SUN RGB-D slice (Song, Lichtenberg, and Xiao 2015), retaining the ten furniture-centric categories chair, table, desk, sofa, bed, nightstand, bookcase, dresser, toilet, bathtub. To assess cross-domain generalization we curate additional 10-class split: an IN10 set that merges ARKitScenes (Baruch et al. 2021) and Hypersim (Roberts et al. 2021). Following Cube R-CNN, we discard objects with invalid depth or marginal image overlap and fix a handful of boxes whose third rotation-matrix column is not aligned with gravity.
We report seven metrics taken from Cube R-CNN and NOCSFormer (Brazil et al. 2023; Krishnan et al. 2024): $IoU_{3D}$, TRANS [cm], ROT [deg], SIZE [cm], 5GRAV, 10HEAD and $mAP_{0.5}$. $IoU_{3D}$ is computed with the exact volumetric intersection between predicted and ground-truth cuboids, not the fast approximation used in the original Cube R-CNN code. AP denotes mean Average Precision at a fixed 2D IoU threshold of 0.5. $mAP_{0.5}$ averages AP over the nine 3D IoU thresholds $\{0.05,0.10,\dots,0.50\}$, providing a finer-grained view of performance. 5GRAV is the percentage of predictions with gravity-axis error $5^{\circ}$; 10HEAD is the percentage with heading (yaw) error $10^{\circ}$ (Krishnan et al. 2024). For $IoU_{3D}$, 5GRAV, 10HEAD and $mAP_{0.5}$ higher values indicate better performance, whereas lower values are preferred for TRANS, ROT, and SIZE.
@@
SUN10 43.21 11.60 22.91 12.93 78.03 47.32 35.70
SUN10 + 3DRot 44.51 11.19 20.93 12.45 79.89 55.53 38.11
IN10 30.39 22.97 31.34 13.62 79.71 38.01
IN10 + 3DRot 30.59 22.75 31.23 13.57 81.20 37.80
@@
BS (Baseline) 43.21 11.60 22.91 12.93 78.03 47.32 35.70
BS + 0.5 F 43.35 11.24 25.36 12.68 26.59 47.97 35.84
BS + 0.5 F(kc) 43.79 11.43 20.91 12.61 78.67 55.23 36.58
BS + 0.8 R10-5-5 44.43 11.10 21.16 12.79 78.94 54.17 38.29
BS + 0.8 R30-5-5 43.89 11.22 21.70 12.60 81.30 52.38 37.53
BS + 0.8 R10-5-5 + 0.5F(kc) 44.51 11.19 20.93 12.45 79.89 55.53 38.11
Table 2: Ablation study on SUN10: effect of each augmentation component. All runs use the same training recipe as Table 1.
Method Train set $IoU_{3D}\,\uparrow$
Cube R-CNN SUN RGB-D 36.2
Cube R-CNN OMNI3DIN 37.8
Ours SUN RGB-D 43.21
Ours + 3DRot SUN RGB-D 44.51
Table 3: Evaluated on the SUN RGB-D 10-category split, using a category set closely aligned with Cube R-CNN
Method Center Chir. $IoU_{3D}\,\uparrow$ ROT $\downarrow$ $mAP_{0.5}\,\uparrow$
Baseline 43.70 22.22
3DRot (full) $\surd$ $\surd$ 44.58 20.78 38.06
w/o Center $\times$ $\surd$ 44.43 20.81 38.06
w/o Chirality $\surd$ $\times$ 44.40 23.57 37.49
Table 4: Ablation on SUN10. To avoid information loss caused by the model’s downsampling, objects whose height is less than 6.25% of the image height are removed. Center: image padding & principal-point realignment. If Center is disabled, we simply crop to the rotated image’s post‑rotation width and height; Chir.: chirality preservation during flips.
Implementation Details
We adopt DINO-X (Ren et al. 2024; Liu et al. 2024), a DETR-style detector whose hybrid CNN-Transformer encoder produces highly accurate 2D boxes and class logits across diverse indoor datasets. Although the original model is pre-trained with open-vocabulary text supervision, we use it purely as an off-the-shelf 2D object recognizer: all backbone and detection layers are kept frozen during our experiments.
The last decoder layer produces N=900N{=}900 object queries (256 dims). We feed the corresponding image features into a Cube R-CNN head (Brazil et al. 2023), which predicts a 9-D cuboid representation-2D center, depth, 3D dimensions, and a full $3{\times}3$ rotation matrix.
Main Results
Table 1 summarizes single-run performance (seed 42) on the two evaluation splits. The first two rows correspond to the SUN10 setting: a vanilla DINO-X + Cube R-CNN baseline (labeled SUN10) and the same backbone equipped with 3DRot. Unless stated otherwise, 3DRot is applied with probability 0.8 for an in-plane rotation (yaw $\pm 10^{\circ}$, pitch/roll $\pm 5^{\circ}$) and probability 0.5 for a chirality-preserving horizontal flip; flips may be composed with the rotation within the same image.
On SUN10, 3DRot improves every metric: $IoU_{3D}$ rises by 1.30, translation and size errors fall by 3.5% and 3.7%, and rotation error decreases by 8.6%. The largest relative gain is on orientation accuracy: 10HEAD increases from 47.3% to 55.5%, which also raises $mAP_{0.5}$ from 35.7% to 38.1%. For the IN10 indoor split the same augmentation brings a smaller yet consistent benefit ($IoU_{3D}$ +0.20, translation -0.22 cm), showing that 3DRot does not overfit to SUN10 statistics.
Table 3 offers an additional perspective. All methods are evaluated on the identical SUN RGB-D test split; minor differences in training-set preprocessing do not affect the comparability of the reported results. The two Cube R-CNN lines are taken from the original Cube R-CNN paper: one model is trained only on SUN RGB-D, the other on the larger OMNI3DIN aggregate (SUN RGB-D, ARKitScenes, Objectron and Hypersim). Our detector, trained solely on SUN RGB-D and augmented with 3DRot, gains 1.30 $IoU_{3D}$ over its own SUN baseline, providing a performance increase similar to that achieved by Cube R-CNN when switching to OMNI3DIN, yet without any extra training images.
Ablation Study
Notation.
Baseline denotes the vanilla DINO-X + Cube R-CNN pipeline without any geometric augmentation. 0.5 F applies a horizontal flip with probability 0.5. (kc) indicates keep chirality: the mirrored image plane is mapped back to a right-handed camera frame. 0.8 R10-5-5 and 0.8 R30-5-5 sample yaw, pitch and roll from the symmetric angle ranges (in degrees) and apply them with probability 0.8. Flip and rotation are composed whenever both Bernoulli trials succeed.
Individual operators (Table 2).
Adding flips alone (BS + 0.5 F) slightly raises $IoU_{3D}$ but increases rotation error to 25.36° and collapses gravity alignment, indicating that a plain reflection is harmful. Activating keep-chirality (BS + 0.5 F(kc)) restores a low rotation error (20.91°) and yields a 10HEAD gain of 7.9. Pure rotation improves most metrics; the narrow band R10-5-5 attains the highest AP, whereas the wider yaw range R30-5-5 excels on 5GRAV. The full recipe-R10-5-5 combined with chirality-safe flipping-achieves the best overall balance: lowest size error (12.45 cm), highest 10HEAD (55.53) and competitive detection scores ($mAP_{0.5}$ = 38.11).
Center vs. Chirality (Table 4).
Padding the rotated image and realigning the principal point (column Center) brings a small but consistent gain: removing it degrades $IoU_{3D}$ from 44.58 to 44.43 and leaves $mAP_{0.5}$ unchanged. Chirality preservation is critical: turning it off inflates the rotation error from 20.78° to 23.57° and lowers $mAP_{0.5}$ by 0.57, confirming that naively mirroring a full $3\times 3$ rotation matrix breaks orientation consistency.
Conclusion & Discussion
We have introduced 3DRot, a simple yet effective rotation-and-reflection augmentation that operates directly at the camera optical center. 3DRot does not require depth information for scene reconstruction while maintaining 2D-3D geometric consistency. Extensive experiments on SUN10 and IN10 show that 3DRot improves $IoU_{3D}$, rotation, gravity alignment and $mAP_{0.5}$ without any change to network architecture or loss weights. The gains generalize across domains and, on SUN RGB-D, rival those obtained by training on a dataset an order of magnitude larger.
Because 3DRot operates purely on camera extrinsics and a closed-form homography, it is readily transferable to other 3D vision tasks whose labels can be rotated in tandem, such as category-level NOCS alignment, human/robot keypoint estimation, or dense depth completion. Similarly, LiDAR-centric detectors could benefit from a hybrid variant that rotates point clouds and images jointly (Wang et al. 2021; Zhang, Wang, and Loy 2020).
Most indoor and outdoor benchmarks are gravity-aligned and near-vertical. The cameras’ roll angles during capture are tightly concentrated around zero. Rotation-centric augmentation is therefore especially valuable for domains with extreme or rapidly changing camera poses-e.g. wrist-mounted industrial manipulators or aerial drones-where robustness to arbitrary pitch, roll and yaw is critical.
All geometry-consistent augmentations-including 3DRot-can also serve as annotation auditors. Training two otherwise identical models, one with and one without such an augmentation, and then comparing their results: a well-annotated dataset yields similar or higher performance by using geometry-consistent augmentations.

### v3

@@ Appendix E Image Padding and Principal-Point Realignment
Appendix F Depth Metrics
For completeness, we recall the standard single-image depth metrics from Eigen et al. (Eigen et al., 2014). Let $d_{i}$ and $\hat{d}_{i}$ denote the ground-truth and predicted depth at pixel $i$, and $N$ the number of valid pixels. We report absolute relative error (abs-rel), root mean squared error (rmse), and the $\delta<1.25$ accuracy (a1), defined as
$\text{abs-rel}=\frac{1}{N}\sum_{i}\frac{|d_{i}-\hat{d}_{i}|}{d_{i}}$
$\text{rmse}=\sqrt{\frac{1}{N}\sum_{i}(d_{i}-\hat{d}_{i})^{2}}$
$\delta<1.25=\frac{1}{N}\bigl|\{i:\max(d_{i}/\hat{d}_{i},\hat{d}_{i}/d_{i})<1.25\}\bigr|$
Appendix G Implementation Details
G.1 Monocular 3D detection (SUN10/IN10)
The last decoder layer produces N=900N{=}900 object queries (256 dims). We feed the corresponding image features into a Cube R-CNN head (Brazil et al., 2023), which predicts a 9-D cuboid representation-2D center, depth, 3D dimensions, and a full $3{\times}3$ rotation matrix.
G.2 Monocular depth estimation (NYU Depth v2 / SUN RGB-D)
For depth estimation we use the BTS ResNet-50 model from Monocular-Depth-Estimation-Toolbox, implemented as a DepthEncoderDecoder with a BTSHead and the standard scale-invariant loss. On NYU Depth v2 we train for 60 epochs with AdamW (learning rate $1{\times}10^{-4}$) and a OneCycle learning-rate schedule, using a batch size of 64 and $416{\times}544$ crops after the NYU-specific NYUCrop. 3DRot is an operator inserted after NYUCrop and before the usual RandomFlip, RandomCrop, and ColorAug: with probability 0.7 it applies a small camera-centric rotation (yaw/pitch up to $3^{\circ}$, roll up to $5^{\circ}$) and updates the RGB image, depth map, and camera intrinsics consistently. All other data-processing steps and evaluation metrics follow the official NYU Depth v2 configuration.
G.3 LiDAR+RGB 3D detection (KITTI / MVX-Net)
For LiDAR+RGB detection we follow the official MVX-Faster-RCNN configuration in MMDetection3D, using a DynamicMVXFasterRCNN detector with a ResNet-50–FPN image backbone, a DynamicVFE + SparseEncoder LiDAR branch, and an Anchor3DHead. We train on the KITTI three-class benchmark for 50 epochs with AdamW (learning rate $3{\times}10^{-3}$), a cosine-annealing schedule with linear warm-up, and a batch size of 8, starting from the released pre-trained MVX-Faster-RCNN checkpoint. The baseline pipeline includes GlobalRotScaleTrans and RandomFlip3D as standard 3D scene-level augmentations. 3DRot is implemented as an 3DRot operator placed before these two transforms: with probability 0.3 it samples a camera-centric rotation with yaw and pitch in $\pm 3^{\circ}$ (zero roll) and applies it jointly to the RGB image, calibration matrices, LiDAR points, and 3D boxes. All other hyperparameters and evaluation metrics are kept identical to the default MVX-Net KITTI setup.
Appendix H Impact of Approximate Intrinsics
This section elaborates on how approximate or heterogeneous camera intrinsics affect our augmentation and how they can be handled in practice.
Estimating intrinsics from 2D–3D correspondences.
In many 3D vision benchmarks, camera intrinsics are either provided or can be reasonably estimated. Whenever a task supplies 2D pixel coordinates together with corresponding 3D points, one can recover the full camera matrix from a small set of 2D–3D correspondences. The camera parameters—intrinsics, rotation, and translation—jointly have 11 degrees of freedom. Each 2D–3D correspondence contributes two independent constraints, because the projection of a 3D point carries its own unknown depth/scale. Therefore at least six non-degenerate correspondences are required ($2\times 6\geq 11$). If these six 3D points are non-coplanar, the constraints are independent and one can solve for the full camera matrix $P\in\mathbb{R}^{3\times 4}$ and then decompose it as
$P=K\,[R\mid t],$
where $K$ is upper-triangular with positive diagonal entries and $(R,t)$ are the extrinsics.
Effect of approximate intrinsics.
When the estimated intrinsics $K$ are only approximate, the image warp and label update in 3DRot may become slightly mismatched. Empirically, moderate perturbations in $K$ tend to behave like a smooth re-parameterization of the camera rays: the mapping from pixels to rays in $\mathbb{R}^{3}$ is distorted but remains consistent across the image. Modern detectors are usually robust to such structured perturbations and can absorb them during training, so the gains from 3DRot typically degrade gracefully rather than collapsing.
Normalizing heterogeneous intrinsics via a virtual camera.
A related concern is that large diversity in intrinsics across a dataset might limit the applicability of camera-centric augmentations. In practice, this diversity can be absorbed by adopting a shared virtual camera parameterization as in Cube R-CNN (Brazil et al., 2023). Concretely, consider a real camera with intrinsics $K$ (focal length $f$) and image height $H$, and a fixed virtual camera $(K_{v},f_{v},H_{v})$. For a 3D point with metric depth $z$ in the real camera, the corresponding virtual depth $z_{v}$ is
$z_{v}=z\cdot\frac{f_{v}}{f}\cdot\frac{H}{H_{v}}.$
This change of variables is a similarity transform in projective space. A 3D point $\mathbf{X}$ that satisfies
$s\,\mathbf{x}=K\,[R\mid t]\,\mathbf{X}$
in the real camera can be equivalently represented in the virtual camera as
$s_{v}\,\mathbf{x}=K_{v}\,[R_{v}\mid t_{v}]\,\mathbf{X}_{v},$
where $\mathbf{X}_{v}$ encodes the same scene geometry but uses $z_{v}$ instead of $z$. Our augmentation only requires that 3D points and 2D pixels obey a consistent pinhole projection model. Therefore, 3DRot can be applied directly in this unified virtual-camera space: (i) normalize all annotations to the virtual camera, (ii) perform geometry-consistent 3D rotations and reflections there, and (iii) optionally map the augmented samples back to the original cameras if needed.
In summary, when exact intrinsics are unavailable, one can either estimate them from sparse 2D–3D correspondences or absorb their variability into a virtual-camera parameterization. In both cases, 3DRot remains valid as long as the chosen camera model defines a consistent mapping between image pixels and 3D rays.
Appendix I Additional Quantitative Results
Additional quantitative results are shown in Tables 9, 10, 11, and 12.
Table 9: Validation results of DINO-X and Cube R-CNN head on the SUN RGB-D 10-category split with different augmentations. Trans and Size are reported in centimeters; Rot is the mean absolute rotation error in degrees. 5GRAV denotes accuracy within $5^{\circ}$ of the ground-truth gravity axis, and 10HEAD denotes accuracy within $10^{\circ}$ of the ground-truth heading. All configurations are trained without any flipping augmentation.
@@ Appendix I Additional Quantitative Results
BASE 42.43 11.87 23.99 13.09 77.03 45.39
BASE + Color jitter 41.51 12.24 22.48 13.07 78.84 48.16
BASE + 3DRot 44.41 11.13 21.91 12.50 77.26 49.06
Table 10: BTS ResNet-50 depth estimation on NYU Depth v2 and SUN RGB-D under different rotation schedules.
@@
No AUG 0.1841 0.6088 0.7089 0.2449 0.6676 0.6065
2DROT 2.5 (baseline) 0.1783 0.5683 0.74722 0.2502 0.6738 0.6049
2DROT 5 0.1775 0.5707 0.74602 0.2477 0.6717 0.6047
2DROT 10 0.1729 0.5704 0.74793 0.2460 0.6776 0.5996
3DROT 1-1-2.5 0.1711 0.5679 0.75202 0.2328 0.6643 0.6138
3DROT 1-1-5 0.1711 0.5714 0.7496 0.2358 0.6607 0.6178
3DROT 2-2-5 0.1685 0.5670 0.7548 0.2333 0.6612 0.612
3DROT 3-3-5 0.1683 0.5695 0.7536 0.2352 0.6633 0.6117
3DROT 2-2-10 0.1741 0.5829 0.7430 0.2375 0.6675 0.6086
Table 11: Full BTS ResNet-50 results with all augmentation variants. Models are trained on NYU Depth v2 and evaluated on NYU Depth v2 and SUN RGB-D. base uses the default BTS schedule (including 2D rotation and horizontal flips); proj_family applies projective augmentations (flip, in-plane 2D rotation baseline, or our camera-centric 3DRot2-2-5/3DRot3-3-5); color_family adds stronger photometric jitter.
NYU Depth v2 SUN RGB-D
[…]
base plain 0.2105 0.6765 0.6582 0.2891 0.7587 0.5366
proj_family flip 0.2027 0.6828 0.6574 0.2704 0.7408 0.5485
proj_family 2DRot 0.2059 0.6460 0.6818 0.2835 0.7567 0.5343
proj_family 3DRot2-2-5 0.1940 0.6370 0.6954 0.2684 0.7414 0.5494
proj_family 3DRot3-3-5 0.1965 0.6430 0.6918 0.2733 0.7504 0.5429
color_family color 0.1879 0.6102 0.7134 0.2532 0.6867 0.5932
Table 12: Full MVX-Net ablations on KITTI under different 3DRot settings (yaw–pitch–roll–probability). We report moderate AP3D, APBEV, and AOS for Overall and each class.
Overall Pedestrian Cyclist Car
Exp. 3D BEV AOS 3D BEV AOS 3D BEV AOS 3D BEV AOS
BASE 63.85 70.79 66.60 55.12 62.90 45.25 58.13 61.48 62.01 78.31 88.00 92.53
3-0-0-0.3 62.97 69.63 66.76 56.25 63.78 47.74 56.12 59.38 60.52 76.55 85.74 92.02
0-3-0-0.3 64.87 71.54 68.96 57.78 64.32 49.86 59.32 63.17 64.50 77.52 87.13 92.51
0-0-3-0.3 58.58 63.84 53.43 54.98 59.49 30.92 45.56 45.98 38.84 75.21 86.06 90.52
3-3-3-0.3 59.38 67.11 62.62 47.18 56.20 41.95 52.90 57.89 53.71 78.07 87.25 92.19
1-1-0-0.3 62.78 69.68 67.31 53.30 60.97 50.35 56.27 59.73 59.47 78.77 88.34 92.10
3-3-0-0.3 65.16 71.09 67.18 59.26 66.40 46.87 57.94 59.35 62.44 78.27 87.50 92.23
5-5-0-0.3 63.40 70.75 67.56 55.54 65.44 45.50 53.71 58.09 60.49 79.20 88.19 92.03
3-3-0-0.1 63.50 70.64 65.53 54.02 62.39 40.94 57.57 60.57 60.56 78.74 87.92 92.45
3-3-0-0.5 63.47 69.71 67.02 56.43 63.27 47.90 55.29 58.00 60.86 78.70 87.85 92.29